SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report June 27, 2003 Day 178

Status Report From: National Oceanic and Atmospheric Administration
Posted: Friday, June 27, 2003

DMSP

All DMSP operations were nominal over the past 24 hours.

POES

NOAA-12 rev 62945 / F at 1330z on June 26:  Executed command procedure as per RCP #12_106_1 to change Solar Array offset from 0 degrees to -30  degrees.  Historical data confirmed that the -30 degree array offset occurred at approximately 1402z after entering eclipse (note:  NOAA-12 flight software is different than the other s/c; on the other s/c, the array slew would have been immediate) .  At 1510z, command SD1WD 09B5 was sent as per RCP# 12_108_1 to assess the effects of previous array slew commanding.

All POES operations were nominal over the past 24 hours.

GOES

During the GOES GIMTACS software release that started at at 1230z on June 26, complications with a COMM controller configuration caused a temporary loss of command which caused the GOES-10 1245Z PACUS image, and GOES-12 1245Z Northern Hemi image and 1246Z sounding to be missed.  By yesterday evening, all command and control software release issues were resolved and tested.

At 1451z on June 26, the GOES-11 DCPR frequency was set to the domestic frequency (401.9 MHz) as per SCR to accommodate WCDAS testing.

GOES-11 ranging with Goldstone, Santiago and NASA Wallops continued throughout yesterday.  An early analysis of the orbit indicates that an E/W maneuver (to offset residual drift from the N/S maneuver) will be required and is expected to occur on Monday.

INR engineers are still trying to resolve a GOES-10 E/W image shift noted early this morning.  It is believed that this shift is independent of the GIMTACS software release.  Engineers discovered that the SPS database defined subsat point had changed from 135.5 to 135.0 this morning at WCDAS.  The correct value was reconfigured this AM and engineers are reviewing the event to understand how the database value was changed.

At 0706z on June 27, a GOES-12 imager frame break was observed and attributed to loss of signal synchronization.

At 0914z on June 27, a GOES-9 imager frame break was observed and attributed to loss of signal synchronization.

GOES-8 continues its westward drift at 1.06 deg/day. The drift table with the estimated current position can be found at http://www.oso.noaa.gov/goes/GOES-8_Drift_Table.txt

 

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.